Determine the net capacitance C of each network of capacitors shown below. Assume that =1.0pF, =2.0pF,=4.0pF and =5.0pF. Find the charge on each capacitor, assuming there is a potential difference of 12.0V across each network.



1
Expert's answer
2022-02-10T08:37:07-0500

Answer

All charges in parallel so

Equivalent capacitance is

Given by

"C_e=C_1+C_2+C_3+C_4\\=1+2+4+5\\=12pF"

V=12V

Charge on each capacitor

"Q_1=C_1V=1pF\times12=12pC\\Q_2=C_2V=2pF\times12=24pC\\Q_3=C_3V=4pF\times12=48pC\\Q_4=C_4 V=5pF\times12=60pC"